OpenStudy (haleyelizabeth2017):

graph the system of constraints and find the value of x and y that maximize the objective function: x>=0 y>=0 y<=3 y<=-x+5 C=-5x+3y answer options in reply

OpenStudy (haleyelizabeth2017):

a. (0, 0) b. (2, 3) c. (5, 0) d. (0, 3)

OpenStudy (haleyelizabeth2017):

@micahm this is what I need help with and 2 others :(

OpenStudy (haleyelizabeth2017):

@TheSmartOne @Haichi

OpenStudy (haleyelizabeth2017):

I don't get this :'(

TheSmartOne (thesmartone):

@aum @jdoe0001 can probably help you

OpenStudy (haleyelizabeth2017):

Thank you

OpenStudy (aum):

Draw each inequality on the same graph. Find the solution space. Evaluate the function at each vertex on the boundary and find which one maximizes C.

OpenStudy (haleyelizabeth2017):

That's still confusing....I'm normally good at math but my brain isn't working and I still have a Gov test to do :'(

OpenStudy (haleyelizabeth2017):

one second. I will try it

OpenStudy (aum):

Skip the first two inequalities. All it means is the first quadrant.

OpenStudy (haleyelizabeth2017):

That's even more confusing

OpenStudy (aum):

y >= 0 means above and including the x-axis x >=0 means to the right and including the y-axis. That is the first quadrant with both x and y axes included.

OpenStudy (haleyelizabeth2017):

Okay :) thank you......one second....almost done

OpenStudy (aum):



OpenStudy (aum):

Find C at (0,0); (5,0); (2,3); (0,3) and see which one maximizes C.

OpenStudy (haleyelizabeth2017):

how? I'm sorry to be taking so much time....this is just really confusing for me and I have no idea why :'( so like plug in x and y and see which on like is greater than c or something?

OpenStudy (haleyelizabeth2017):

I have no idea

OpenStudy (aum):

The first four inequalities are the constraints. You can select x and y values only from the region that satisfies all 4 inequalities. In the diagram above, the region marked "Solution Space" is the region that satisfies all 4 inequalities. The points of interest are those that lie on the boundary. Test each point and see which one gives the largest C. C = -5x + 3y at (0,0); C = 0 at (5,0); C = -25 at (2,3); C = -10 + 9 = -1 at (0,3); C = 0 + 9 = 0 C is maximum at (0,3)

OpenStudy (haleyelizabeth2017):

Thank you so much! OH!!!! so it it the most like um......biggest amount?

OpenStudy (haleyelizabeth2017):

THANK YOU SO MUCH FOR HELPING ME @aum!

OpenStudy (aum):

Yes. You are welcome.

OpenStudy (haleyelizabeth2017):

:) I got it now

OpenStudy (aum):

Good.